OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 (3d)3(d) Expression #2 (3d2)2

OpenStudy (anonymous):

The anwsers are The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

use the commutative property of multiplication 3d3d = (3*3)(d*d)

OpenStudy (anonymous):

same with the second expression (2*3*2)d

OpenStudy (anonymous):

so would it be the second one. The exponents of expression 1 are greater than expression 2, right?

OpenStudy (anonymous):

whats 3*3= d*d=

OpenStudy (anonymous):

wouldnt it be 6d^2 ?

OpenStudy (anonymous):

unless you left an addition sign out of the original express it should be 9d^2. 3 times 3 is 9

OpenStudy (anonymous):

expression*

OpenStudy (anonymous):

the second expression is 2 times 2 times 3=

OpenStudy (anonymous):

4 times 3 is?

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

so you have 9d^2 and 12d. Which has the greater exponent

OpenStudy (anonymous):

OH, OK

OpenStudy (anonymous):

so the exponents would be the same for the expressions

OpenStudy (mathmale):

This problem has become harder, a lot harder, than it needs to be, because copying and pasting the original problems has not expressed exponentiation properly. Either use Equation Editor (below) or the Draw utility to present the original problem. Or, use the symbol " ^ " to express exponentiation. (3d2)2 is subject to misinterpretation. Is that final 2 an exponent or a multiplicand? Here is one possible interpretation:\[(3d^2)^2\]

OpenStudy (mathmale):

I am not saying this is correct. But I do need for you to express your math symbols in such a way that it is not necessary to ask you, "Do you mean ... this ... or do you mean ... that." Please re-write the given expressions using any of the 3 methods mentioned above. Then we can finally answer the question with confidence.